OpenStudy (anonymous):

A purchase is made under the following conditions. There is a 25% discount due to a sale, another 10% discount because of the size of the purchase, and a 3% discount for paying cash. These discounts are successive. That is, they are taken one after the other. The customer paid $195.89. What was the original purchase price before any discounts were considered?

OpenStudy (help!!!!):

so for this problem u would the opposite of the normal way

OpenStudy (anonymous):

so what is the process?

OpenStudy (help!!!!):

195.89-(195.89x.10)

OpenStudy (help!!!!):

so after that u put out your answer_______-(_____x.03

OpenStudy (anonymous):

what about the 25%

OpenStudy (help!!!!):

you would do the same

OpenStudy (help!!!!):

______-(____x.25)

OpenStudy (anonymous):

so would you do the 25 percent before the 10% or after the 3% ??

OpenStudy (help!!!!):

:D

OpenStudy (help!!!!):

it really doesnt matter thought i would do it the way the problem is in order

OpenStudy (anonymous):

ok so after putting in the answers and getting the final one, is that the original price?

OpenStudy (help!!!!):

so i would have done 25,10,3

OpenStudy (help!!!!):

yes ur CORRect :D!!!!!!!

OpenStudy (help!!!!):

an medal would be appreciated

OpenStudy (anonymous):

ok thank you so much!

OpenStudy (anonymous):

?=195.89*1.25*1.1*1.03=277.42921

OpenStudy (mathstudent55):

Here's one way of solving this. Let's say you start with a $100 item and you do the three discounts in series: $100 item minus 25% discount = 75% x $100 = $75.00 $75 item minus 10% discount = 90% x $75 = $67.50 $67.50 item minus 3% discount = 97% x $67.50 = $65.475 This means that a $100 item ends up costing $65.475, or 65.475% of its original price. Now we know that an item that undergoes these three discounts in series ends up costing 65.475% of the original price. This means that $195.89 is 65.475% of the original price. So the quesion now is $195.89 is 65.475% of what? Use a proportion: $195.89 : 65.475% = x : 100% x = 100% x $195.89 / 65.475% x = $299.18 The original price was $299.18